### v2

# Toward a unified theory for common affine roots of general sets of multivariate polynomials

Abstract
For univariate polynomials over arbitrary field the degree gives an upper bound on the number of roots (factor theorem) and as a related result for any finite point-set one can construct a polynomial of degree equal to the cardinality having all the points as roots (interpolation theorem). Terence Tao noted in [39] that the theory of multivariate polynomials is not yet sufficiently matured to provide similar theorems with an equally simple relation between them. In the present paper we argue that for general multivariate polynomials the right measure for the size of the polynomial should not be the degree, but the leading monomial. In this setting the footprint bound [23] becomes a natural enhancement of the factor theorem providing a bound on the number of common roots of general multivariate polynomials which is sharp for all finite Cartesian product point sets. Furthermore, by using methods from the theory of error-correcting codes we establish a natural formulation of the interpolation theorem to the case of common roots of multivariate polynomials. In short the two theorems reduce to the same result, but for dual spaces, establishing the unification requested in [39]. We leave it for further research to possibly establish similar related results taking multiplicity into account.
Keywords: Affine roots, footprint bound, generalized Alon-Füredi bound, interpolation, multivariate polynomials
MSC: 05E40; 12E05
1 Introduction
For a univariate non-zero polynomial over an arbitrary field the number of roots is at most equal to its degree, and if the field is algebraically closed and if roots are counted with multiplicity then equality holds. This result is in close correspondence with the fact that the interpolation problem of finding a polynomial of lowest possible degree evaluating to 0 on a finite point-set boils down to calculating the product of linear terms each having one of the desired elements as root, and by repeating terms one can also deal with multiplicity. The situation is very much in contrast to that of multivariate polynomials. Firstly, a multivariate polynomial ring is not a unique factorization domain, and therefore the considered interpolation problem is not about multiplying linear factors. Secondly, a non-constant multivariate polynomial over an infinite field has infinitely many roots, and therefore if the aim is to consider a case of finitely many roots then one either needs to restrict to roots from a prescribed finite point-set or one must study mutual roots of a set of polynomials which together define a zero-dimensional ideal.
The purpose of the present paper is to establish a natural enhancement of the results from univariate polynomials into the case of general multivariate polynomials and by doing so to establish a connection between the following two problems:
  • (I)
    How many common roots can a finite set of polynomials have given some measure on their sizes?
  • (II)
    Given some measure on the sizes of polynomials let $a,k$ be any positive integers. If we are to guarantee that for any point-set $A$ of cardinality at most $a$ there exist $k$ polynomials with pairwise different leading monomials having $A$ as common roots, what is then the sharpest upper bound in terms of the measure on sizes of polynomials we can formulate?
Our first step in this direction is to replace in our analysis the concept of degree with that of leading monomial being then the chosen measure in the present paper. For univariate polynomials there is only one choice of monomial ordering and the leading monomial of a polynomial $F(X)$ of degree $d\geq 0$ equals $X^{d}$, hence the two concepts are basically the same. In contrast, for multivariate polynomials there are infinitely many monomial orderings and for many of these the total degree of the leading monomial does not even need to be equal to the total degree of the polynomial under consideration. Considering leading monomial(s) is not only as natural a choice for our study as considering total degree(s); as shall be demonstrated it is the key to achieve the above mentioned enhancement and connection. Even more importantly, using the leading monomial rather than the total degree provides much more fine-grained information which is sharp in more cases.
In his exposition [39] Terence Tao treats problems (I) and (II) from the above paragraph including discussing on the one side Alon’s famous combinatorial Nullstellensatz [1], the Schwartz-Zippel bound [42, 36, 29] (as well as Dvir’s generalization taking multiplicity into account [13, 14, 15]) and on the other side the below standard result [39][Lem. 1] on interpolation; with the aim of solving problems in extremal combinatorics:
Proposition 1.
Let $\mathbb{F}$} be a field, let $m\geq 1$ be an integer, and $d\geq 0$. If $A\subset{\mathbb{F}}^{m}$ has cardinality less than ${d+m}\choose{m}$, then there exists a non-zero polynomial $F\in{\mathbb{F}}[X_{1},\ldots,X_{n}]$ of degree at most $d$ having $A$ as roots.
At pages 27-28 Tao writes:
Unfortunately, the two methods cannot currently be easily combined, because the polynomials produced by interpolation methods are not explicit enough that individual coefficients can be easily computed, but it is conceivable that some useful unification of the two methods could appear in the future.
The theory of commutative algebra and algebraic geometry is very rich and indeed Tao in addition to the above results treats a cornucopia of deep results such as the Alon-Füredi bound, Bezout’s Theorem, Stepanov’s Method, the Hasse Bound and many more [4, 8, 27, 40, 30, 37, 41], as well as provide interpolation counterparts in more cases. Our contribution is to address the problem of common roots of general polynomials where the only assumption is that no two of them have the same leading monomial which is no restriction as we can always apply the division algorithm [10][Sec, 2, Par. 3]] without effecting the set of common roots. Similarly, we consider any finite point-sets without any restriction on their structure. Therefore, our answer to (II) should be compared to Proposition 1 and our answer to (I) should be compared to the Schwartz-Zippel bound, the Alon-Füredi bound and the generalized Alon-Füredi bound [7]. In both cases the material of the present paper can be viewed as improvements and of course also as enhancements as we do not restrict to single polynomials but rather consider arbitrary finite set of polynomials.
In the present paper we give refined answers to both (I) and (II) taking the leading monomial into account and we enhance the methods to treat the case of common roots of a set of more polynomials. Knowing the leading monomial corresponds to having important information regarding the most crucial coefficient and thereby our insight addresses the remark by Tao. More importantly, we demonstrate a very strong relationship between our answers to (I) and (II) which is then our proposal for the requested useful unification, except we do not treat multiplicity. We leave it for further research to explore implications in the theory of extremal combinatorics.
At the heart of our method is insight from the theory of error-correcting codes which we translate to hold for general fields rather than just for finite fields. Our answer to (I) is basically to recall the footprintbound [23] in some of its many variants including enhancing a formulation known as the Feng-Rao bound for primary codes [3, 20, 24] to hold for arbitrary field. When applied to finite Cartesian product point-sets the footprint bound is sharp in that it is easy to find a polynomial which is the product of linear terms having exactly as many roots as predicted by the bound. Our answer to (II) is almost trivial in the case of only one monomial, but to treat common roots of more polynomials we shall need some basic lemmas by Forney [19] on linear codes over arbitrary field. The answer to (II) is then about applying an enhanced version of the Feng-Rao bound for dual codes [16, 17, 18, 28, 25], which then also demonstrates the relationship between the two methods for arbitrary fields as well as demonstrate the sharpness of our answer to (II) in the case of finite fields.
The paper is organized as follows. We start in Section 2 by presenting our main finding regarding (II) and by recalling the footprint bound which addresses (I). In that section we also provide a first description of the relationship between these results. To prepare for the proof of the first mentioned result we then in Section 3 treat related problems in the language of error-correcting codes over arbitrary fields. In Section 4 as a corollary we then obtain the proof of our answer to (II) which also demonstrates the strong relationship between (I) and (II), and which implies that our answer to (II) is sharp in the case of finite fields. Section 5 contains concluding remarks.
2 Main results on multivariate polynomials
Our main result is the following enhancement of Proposition 1 which surprisingly is best possible as for finite fields we can show it to be sharp. Observe that if $\mathbb{F}$} is a finite field with $q$ elements, in which case we write ${\mathbb{F}}={\mathbb{F}}_{q}$, and if $q\leq d$ then Proposition 1 does not give us any real information as all points in ${\mathbb{F}}^{n}$ are then trivially roots of each of the polynomials $X_{1}^{q}-X_{1},\ldots,X_{m}^{q}-X_{m}$. The below theorem is prepared to also give non-trivial insight in such cases.
Definition 2.
For multivariate monomials $N,N_{1},\ldots,N_{s}$ write
$\mu(N)$ $:=$ $\#\{M\mid M{\mbox{ is a monomial which divides }}N\}$
$\mu(N_{1},\ldots,N_{s})$ $:=$ $\#\{M\mid M{\mbox{ is a monomial which divides some }}N_{i},i\in\{1,\ldots,s\}\}.$
Definition 3.
In the case of polynomials in $m$ variables over a finite field ${\mathbb{F}}_{q}$ define
$\Box:=\{X_{1}^{i_{1}}\cdots X_{m}^{i_{m}}\mid 0\leq i_{1},\ldots,i_{m}\leq q-1\}.$
Remark 4.
Note, that if we write ${\mathbb{F}}_{q}^{m}=\{P_{1},\ldots,P_{n=q^{m}}\}$ then
$\{(M(P_{1}),\ldots,M(P_{n}))\mid M\in\Box\}$
constitutes a basis for ${\mathbb{F}}_{q}^{n}$ as a vector-space over ${\mathbb{F}}_{q}$. In particular for any polynomial $G$ with support not being a subset of $\Box$ there exists a polynomial $H$ with support being a subset of $\Box$ in such a way that $G$ and $H$ evaluates similarly at ${\mathbb{F}}_{q}^{m}$ and such that ${\mbox{lm}}(H)\preceq{\mbox{lm}}(G)$ holds for any choice of monomial ordering $\prec$. Here, ${\mbox{lm}}()$ denotes the leading monomial. More concretely, $H$ can be obtained from $G$ by dividing the latter with $\{X_{1}^{q}-X_{1},\ldots,X_{m}^{q}-X_{m}\}$. Therefore it is no restricting to consider in the finite field case only polynomials with support in $\Box$.
Theorem 5.
Given a monomial ordering $\prec$ on the set of monomials in $m$ variables let $M_{1}\prec\cdots\prec M_{t}$ be strictly consecutive monomials, i.e. for $i=1,\ldots,t-1$ there does not exist any monomial $L$ with $M_{i}\prec L\prec M_{i+1}$. In the special case ${\mathbb{F}}={\mathbb{F}}_{q}$ we shall assume without loss of generality that $\{M_{1},\ldots,M_{t}\}\subseteq\Box$ and allow for $M_{i}\prec L\prec M_{i+1}$ as long as $L\notin\Box$. Consider $1\leq k\leq t$ then for any $A\subseteq{\mathbb{F}}^{m}$ with
$A<\min\{\mu(M_{i_{1}},\ldots,M_{i_{t-k+1}})\mid 1\leq i_{1}<\cdots<i_{t-k+1}\leq t\}$ (1)
there exists $k$ polynomials $F_{1},\ldots,F_{k}$ with pairwise different leading monomials all belonging to $\{M_{1},\ldots,M_{t}\}$ in such a way that every element of $A$ is a common root. If $\mathbb{F}$} is a finite field then the result is sharp meaning that for some $A$ of cardinality one more such polynomials do not exist. Moreover for finite fields one may assume ${\mbox{Supp}}(F_{i})\subseteq\Box$ for $i=1,\ldots,k$.
Proof.
For $k=1$ the proof of Proposition 1 as presented in [39] trivially applies in the following way. Consider the linear map
${\mbox{ev}}:{\mbox{Span}}_{\mathbb{F}}\{M\mid M{\mbox{ divides some }}M_{i},i\in\{1,\ldots,t\}\}\rightarrow{\mathbb{F}}^{\#A}$
given by ${\mbox{ev}}(F)=(F(P_{1}),\ldots,F(P_{\#A\ }))$ where $A=\{P_{1},\ldots,P_{\#A}\}$. As the preimage is larger than the image there must exist a non-zero $F$ in the preimage evaluating to $\vec{0}$. Now ${\mbox{lm}}(F)$ divides some $M_{i}$ and we write $K=\dfrac{M_{i}}{{\mbox{lm}(F)}}$. But then ${\mbox{lm}}(KF)=M_{i}$ and $KF$ of course evaluates to $\vec{0}$. The proof for larger values of $k$ can be found in Section 4, where also the sharpness in case of a finite field is demonstrated. ∎
Remark 6.
Theorem 5 reduces to the well-known result for univariate polynomials by choosing $\{M_{t=1}=X^{d}\}$ where $d$ is the degree under consideration.
Remark 7.
Observe, that in the case of an infinite field or in the case of a finite field ${\mathbb{F}}_{q}$ with $d<q$ Proposition 1 appears as a corollary to Theorem 5, by choosing a graded monomial ordering and by letting $\{M_{1},\ldots,M_{t}\}$ be all monomials of total degree $d$. For Proposition 1 to cover in the case of a finite field ${\mathbb{F}}_{q}$ general meaningful $d$, i.e. $1\leq d\leq m(q-1)$, we request that the support of $F$ is a subset of $\Box$ and we replace $A<{{d+m}\choose{m}}$ with
$A<\#\{M\in\Box\mid\deg(M)\leq d\}.$ (2)
We note that the right hand side of (2) equals the dimension of the q-ary Reed-Muller code ${\mathcal{RM}}_{q}(d,m)$ [12, Lem. 2.4.6]).
We next address (I) by recalling the footprint bound on the number of common roots of any given set of polynomials over arbitrary field [23]. To the best of our knowledge this is the most general and sharp bound for arbitrary sets of polynomials over arbitrary point-sets. Let $\prec$ be a monomial ordering on the set of monomials in $m$ variables, and consider an ideal $I\subseteq{\mathbb{F}}[X_{1},\ldots,X_{m}]$. The footprint of $I$ with respect to $\prec$ is
${\mbox{\hskip-14.22636pt}}\Delta_{\prec}(I):=$
[…]
The crucial observation is that $\{M+I\mid M\in\Delta_{\prec}(I)\}$ constitutes a basis for ${\mathbb{F}}[X_{1},\ldots,X_{m}]/I$ as a vector-space over $\mathbb{F}$} which is seen by applying the division algorithm.
Theorem 8.
If $\Delta_{\prec}(I)$ is finite then the size of the variety $V_{\mathbb{F}}(I)$ is at most equal to $\Delta_{\prec}(I)$ with equality if $I$ is a radical ideal and $\mathbb{F}$} is algebraically closed.
Proof.
Assume first that $V_{\mathbb{F}}(I)$ is finite and write $V_{\mathbb{F}}(I)=\{\beta_{1},\ldots,\beta_{n}\}$. Consider the evaluation map ${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n}$ given by ${\mbox{ev}}(F+I)=(F(\beta_{1}),\ldots,F(\beta_{n}))$. Due to Hermite interpolation this map is surjective and the first part of the theorem follows. To see the last part we use the fact [10][Ch. 4, Sec. 2, Th. 7] that ${\mathcal{I}}(V_{\mathbb{F}}(I))=I$ whenever $I$ is radical and $\mathbb{F}$} is algebraically closed. Aiming for a contradiction we finally assume that $V_{\mathbb{F}}(I)$ is infinite and choose $\{\beta_{1},\ldots,\beta_{n}\}\subseteq V_{\mathbb{F}}(I)$ with $n>\#\Delta_{\prec}(I)$. But then ev cannot be surjective. ∎
As a corollary we obtain the following result where the assumption on $F_{1},\ldots,F_{s}$ can always be ensured due to the division algorithm.
Corollary 9.
Consider $I$ such that $\Delta_{\prec}(I)$ is finite. Let $F_{1},\ldots,F_{s}\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with $F_{1},\ldots,F_{s}\in{\mbox{Span}}_{\mathbb{F}}\Delta_{\prec}(I)$ satisfying that ${\mbox{lm}}(F_{1})=M_{1},\ldots,{\mbox{lm}}(F_{s})=M_{s}$ are pairwise different. Then the number of common roots from $V_{\mathbb{F}}(I)$ is at most
$\#\{M\in\Delta_{\prec}(I)\mid M{\mbox{ is not divisible by any }}M_{i},i=1,\ldots,s\}$
$=$ $\#\Delta_{\prec}(I)-\sigma(M_{1},\ldots,M_{s})$
where $\sigma(M_{1},\ldots,M_{s}):=\#\{M\in\Delta_{\prec}(I)\mid M{\mbox{ is divisible by some }}M_{i},i\in\{1,\ldots,s\}\}$.
As a special case of Corollary 9 we shall obtain a natural counterpart to Theorem 5 in the case of finite Cartesian product point-sets $A_{1}\times\cdots\times A_{m}$. We shall later argue, that it is in a certain sense not a restriction to consider only such point-sets when formulating the counterpart to Theorem 5 (see Remark 11 below). For finite Cartesian product point-sets the $I$ in Corollary 9 becomes
$I=\left\langle G_{1}=\prod_{b\in A_{1}}(X_{1}-b),\ldots,G_{m}=\prod_{b\in A_{m}}(X_{m}-b)\right\rangle$
from which we easily read of the footprint which is independent on the actual choice of monomial ordering $\prec$
$\Delta_{\prec}(I)=\{M{\mbox{ is a monomial }}\mid\deg_{X_{1}}M<\#A_{1},\ldots,\deg_{X_{m}}M<\#A_{m}\}=:\Diamond.$
Note, that if $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$, then $\Diamond=\Box$ where the latter is described in Definition 3.
Corollary 10.
Let $A_{1}\times\cdots\times A_{m}$ be a finite point-set and consider pairwise different monomials $M_{1},\ldots,M_{t}$ in $\Diamond$. Let $\prec$ be an arbitrary monomial ordering. For general set of $k\leq t$ polynomials $F_{1},\ldots,F_{k}\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with ${\mbox{Supp}}(F_{i})\in\Diamond$, $i=1,\ldots,k$ and leading monomials being pairwise different all belonging to $\{M_{1},\ldots,M_{t}\}$ the number of non-common roots of $F_{1},\ldots,F_{k}$ in $A_{1}\times\cdots\times A_{m}$ is at least
$\min\{\sigma(M_{i_{1}},\ldots,M_{i_{k}})\mid 1\leq i_{1}<\cdots<i_{k}\leq t\}$ (3)
$=$ $\min\{\mu\left(\dfrac{N}{M_{i_{1}}},\ldots,\dfrac{N}{M_{i_{k}}}\right)\mid 1\leq i_{1}<\cdots<i_{k}\leq t\}$
where $N=X_{1}^{\#A_{1}-1}\cdots X_{m}^{\#A_{m}-1}$. This bound is sharp.
Proof.
We only need to demonstrate sharpness. To this end write $a_{i}=\#A_{i}$, $A_{i}=\{b_{1}^{(i)},\ldots,b_{a_{i}}^{(i)}\}$ for $i=1,\ldots,m$. To each monomial $N\in\Diamond$ we associate the polynomial
$H_{N}:=\prod_{i=1}^{m}\prod_{j=1}^{\deg_{X_{i}}N}(X_{i}-b_{j}^{(i)})$
which clearly has leading monomial $N$. Then for any set of monomials $N_{1},\ldots,N_{s}\subseteq\Diamond$ the number of non-common roots of $H_{N_{1}},\ldots,H_{N_{s}}$ is precisely $\sigma(N_{1},\ldots,N_{s})$. ∎
A special case of Corollary 10 is when $A_{1}=\cdots=\#A_{m}=:a$ and there is only one polynomial $F$, the only information of which we have is its total degree $\deg F=d$. Writing $d=(\sum_{i=1}^{v}(a-1))+\ell$ where $0\leq\ell<a-1$ we have
$\min\{\sigma(M)\mid M\in\Delta_{\prec}(I),\deg M=d\}$ $=$ $\sigma(X_{1}^{a-1}\cdots X_{v}^{a-1}X_{v+1}^{\ell})$
$=$ $(a-\ell)a^{m-(v+1)}$
which is a special case of the Alon-Füredi bound [2] which follows from Corollary 10 in a similar manner.
Example 1.
Let ${\mathbb{F}}={\mathbb{F}}_{8}$, i.e. the field with $8$ elements, and consider $\langle X_{1}^{8}-X_{1},X_{2}^{8}-X_{2}\rangle\subseteq{\mathbb{F}}_{8}[X_{1},X_{2}]$. Clearly, $V_{\mathbb{F}_{8}}(I)={\mathbb{F}}_{8}^{2}$. We have $\Delta_{\prec}(I)=\{X_{1}^{i_{1}}X_{2}^{i_{2}}\mid 0\leq i_{1},i_{2}<8\}$ regardless of the choice of monomial ordering $\prec$. This set of course is the set $\Box$ from Definition 3. Assume $M_{1}=X_{1}^{3}X_{2}$, $M_{2}=X_{1}^{2}X_{2}^{2}$, and $M_{3}=X_{1}X_{2}^{3}$ and that $\prec$ is the degree lexicographic ordering with $X_{1}\prec X_{2}$ which implies that indeed $M_{1},M_{2},M_{3}$ are strictly consecutive. We have $\mu(M_{1})=\mu(M_{3})=8$, $\mu(M_{2})=9$, $\mu(M_{1},M_{2})=\mu(M_{2},M_{3})=11$, $\mu(M_{1},M_{3})=12$, and $\mu(M_{1},M_{2},M_{3})=13$. From this we conclude that for each $A\subseteq{\mathbb{F}}_{8}^{2}$ of size at most $7$ there exist polynomials $F_{1},F_{2},F_{3}$, ${\mbox{lm}}(F_{1})=M_{1}$, ${\mbox{lm}}(F_{2})=M_{2}$, ${\mbox{lm}}(F_{3})=M_{3}$ having all elements of $A$ as common roots. We similarly conclude that whenever $A$ is of size at most $10$ then there exist $F_{1},F_{2}$ with ${\mbox{lm}}(F_{1})\neq{\mbox{lm}}(F_{2})$, ${\mbox{lm}}(F_{1}),{\mbox{lm}}(F_{2})\in\{M_{1},M_{2},M_{3}\}$ with all elements of $A$ as common roots. Finally for any $A$ with #A<13\#A<13 there exists $F$ with ${\mbox{lm}}(F)\in\{M_{1},M_{2},M_{3}\}$ such that $F(a)=0$ for all $a\in A$. In the other direction $\sigma(M_{1})=\sigma(M_{3})=35$, $\sigma(M_{2})=36$, $\sigma(M_{1},M_{2})=\sigma(M_{2},M_{3})=42$, $\sigma(M_{1},M_{3})=45$ and $\sigma(M_{1},M_{2},M_{3})=46$ which tells us that any polynomial $F$ with ${\mbox{lm}}(F)\in\{M_{1},M_{3}\}$ has at most 6435=2964-35=29 roots and that if ${\mbox{lm}}(F)=M_{2}$ then it has at most 6436=2864-36=28 roots. If ${\mbox{lm}}(F_{1})=M_{1}$ and ${\mbox{lm}}(F_{2})=M_{2}$ or if ${\mbox{lm}}(F_{1})=M_{2}$ and ${\mbox{lm}}(F_{2})=M_{3}$ then $F_{1}$ and $F_{2}$ have at most 6442=2264-42=22. If ${\mbox{lm}}(F_{1})=M_{1}$ and ${\mbox{lm}}(F_{2})=M_{3}$ the number of common roots is at most 6445=1964-45=19. Finally $F_{1},F_{2},F_{3}$ with ${\mbox{lm}}(F_{1})=M_{1}$, ${\mbox{lm}}(F_{2})=M_{2}$, and ${\mbox{lm}}(F_{3})=M_{3}$ have at most 6446=1864-46=18 roots in common.
The similarity in expression (1) and expression (3) suggests that there is a strong relationship between Theorem 5 and Corollary 10 which we now explain.
Remark 11.
Let $M_{1},\ldots,M_{t}$ be as in Theorem 5. If ${\mathbb{F}}={\mathbb{F}}_{q}$ then let $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$. Otherwise, choose finite sets $A_{1},\ldots,A_{m}\subseteq{\mathbb{F}}$ with $A_{i}\geq\mu(M_{1},\ldots,M_{t})$ for $i=1,\ldots,m$ which we adjust in such a way that $A_{1}\times\cdots\times A_{m}\supseteq A$ holds whenever an $A$ is considered. Let $I=\langle\prod_{b\in A_{1}}(X_{1}-b),\ldots,\prod_{b\in A_{m}}(X_{m}-b)\rangle$. Consider the bijective evaluation map
${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n=\#A_{1}\cdots\#A_{m}}$
given by ${\mbox{ev}}(F+I)=(F(P_{1}),\ldots,F(P_{n}))$ where $\{P_{1},\ldots,P_{n}\}=A_{1}\times\cdots\times A_{m}$. Consider
$C_{2}$ $:=$ ${\mbox{Span}}_{\mathbb{F}}\{{\mbox{ev}}(N+I)\mid N\prec M_{1}\}$
$C_{1}$ $:=$ ${\mbox{Span}}_{\mathbb{F}}\{{\mbox{ev}}(N+I)\mid N\preceq M_{t}\}.$
Clearly, $C_{2}\subseteq C_{1}$, with codimension equal to $t$. For $1\leq k\leq t$ consider the relative generalized Hamming weight [31]
$M_{k}(C_{1},C_{2}):=\min\{\#{\mbox{Supp}}(D)\mid D\subseteq C_{1},D\cap C_{2}=\{\vec{0}\},\dim D=k\}$ (4)
originally introduced for finite fields, but here applied to the case of any field. Clearly, for general $F_{1},\ldots,F_{k}$ as in Corollary 10 the smallest attainable number of non-common roots in $A_{1}\times\cdots\times A_{m}$ equals $M_{k}(C_{1},C_{2})$. In the coming sections we shall demonstrate that if $A\subseteq A_{1}\times\cdots\times A_{m}$ satisfies $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ then there exist $k$ linearly independent polynomials as in Theorem 5 which all evaluates to 0 on $A$, but that for some $A\subseteq A_{1}\times\cdots\times A_{m}$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ this does not hold. Here, the dual space (or null-space) is with respect to the usual inner product. Finally, as we shall demonstrate in the next two sections of the paper the right hand side of (1) serves as an upper bound on $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ and we already know that $M_{k}(C_{1},C_{2})$ equals the right hand side of (3). We conclude that problem (I) and problem (II) are basically identical, but applied to dual spaces. All claims of the present remark follows by inspection of the proof of Theorem 5 in Section 4.
We conclude this section by demonstrating the superiority of the footprint bound over other known bounds on roots of general single mulitvariate polynomials. Among these the generalized Alon-Füredi bound [7] is the sharpest as it has as corollary the Alon-Füredi bound [2] which in turn has as corollaries [7] Alon’s combinatorial Nullstellensatz [1] as well as the Schwartz-Zippel bound [42, 36]. For common roots of sets of more general multivariate polynomials we are not aware of other efficient bounds than the footprint bound.
By inspection the generalized Alon-Füredi bound [7, Th. 1.2] can be reformulated as follows.
Theorem 12.
Consider finite sets $A_{1},\ldots,A_{m}\subseteq{\mathbb{F}}$ and a polynomial $F\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with $d_{i}:=\deg_{X_{i}}F<\#A_{i}$, $i=1,\ldots,m$. Let $d$ be the total degree of $F$. The number of non-roots of $F$ in $A_{1}\times\cdots\times A_{m}$ is at least
$\min\{\sigma(M)\mid M{\mbox{ is a monomial}},\deg M=d,{\mbox{ and }}\deg_{X_{i}}M\leq d_{i},i=1,\ldots,m\}.$ (5)
In [6] it is already noted that the generalized Alon-Füredi bound is a consequence of the footprint bound. We here give some more details including discussing when the two bounds may or may not coincide. Consider a graded monomial ordering, i.e. a monomial ordering $\prec$ satisfying that if $\deg N_{1}<\deg N_{2}$ then necessarily $N_{1}\prec N_{2}$. There are many such orderings, including $m!$ degree lexicographic ordering where in addition to the above property ties are broken by the rule that if $K_{1}$ and $K_{2}$ are of the same degree then $K_{1}\prec K_{2}$ if $K_{1}\prec_{lex}K_{2}$ for the given chosen lexicographic ordering. For a degree monomial ordering for sure the leading monomial ${\mbox{lm}}(F)$ belongs to the set considered in (5) and Theorem 12 follows imediately from Corollary 10.
We now comment on the possible tightness of the generalized Alon-Füredi bound. If $d_{1}+\cdots+d_{m}=d$ then there is only one monomial in the support ${\mbox{Supp}}(F)$ of highest degree, hence the two bounds coincide. Moreover, $F$ is a monomial ordering invariant polynomial [21, Def. 8], i.e. all monomials in the support of $F$ divides the unique monomial of highest degree. As shown in [21, Cor. 9] the only polynomials that attain the footprint bound in such a case are the products of linear terms as described in the proof of Corollary 10 (see also [7, Sec. 3.3]). Imagine next that $d_{1}+\cdots+d_{m}\neq d$, in which case there need to be more than one monomial in the argument of (5). If for some $i$ no monomial $M$ in the support of $F$ satisfies simultaneously $\deg_{X_{i}}M=d_{i}$ and $\deg M=d$ then for a lexicographic ordering with $X_{j}\prec_{lex}X_{i}$ for all $j\neq i$ the footprint bound will produce a strictly tighter result than Theorem 12. Example 2 below illustrates the situation. Assume on the other hand that for all $i$ such a monomial exists, then a nesecary condition for Theorem 12 to produce as sharp a result as the footprint bound is that for each of the degree lexicographic orderings the largest monomial in the argument of (5) is of $\sigma$-value equal to (5) which indeed is a very restrictive assumption on the form of $F$.
Example 2.
Consider the Hermitian polynomial $F(X_{1},X_{2})=X_{1}^{q+1}-X_{2}^{q}-X_{2}\in{\mathbb{F}}_{q^{2}}[X_{1},X_{2}]$ where $q$ is a prime-power. We have $d_{1}=q+1$, $d_{2}=q$ and $d=q+1$, hence the argument of (5) is $\{X_{1}X_{2}^{q},X_{1}^{2}X_{2}^{q-1},\ldots,X_{1}^{q+1}\}$ and Theorem 12 therefore tells us that the number of roots of $F$ in ${\mathbb{F}}_{q^{2}}\times{\mathbb{F}}_{q^{2}}$ is at most $(q+1)q^{2}=q^{3}+q^{2}$. Applying instead the footprint bound with the monomial ordering being the lexicographic ordering with $X_{1}\prec_{lex}X_{2}$ we obtain ${\mbox{lm}}(F)=X_{2}^{q}$ from which we see that $F$ can at most have $q^{3}$ roots. It is well-known that actually $F$ possesses precisely this amount of affine roots.
3 Results in linear algebra setting
As a preperation for providing in Section 4 a proof of Theorem 5 as well as giving the missing details of Remark 11 in the current section we treat problems similar to (I) and (II), but in a much more general context than that of multivariate polynomials. Admittedly, for that purpose it would be enough to consider in this general setting only problem (II), as we already proved Theorem 8, Corollary 9 and Corollary 10, but for completeness and to support possible future applications we shall also treat (I).
Algebraic coding theory has been a very active research area for several decades leading to a great number of rich constructions of codes having desirable properties. At the heart of the work is the question of estimating the number of roots of polynomial-like functions in various ${\mathbb{F}}_{q}$-algebras. One attempt to unify some of the theory is the Feng-Rao theory for the ${\mathbb{F}}_{q}$-algebra $({\mathbb{F}}_{q}^{n},+,\ast)$ where $\ast$ is the component-wise (or element-wise) product given by
$(c_{1},\ldots,c_{n})\ast(d_{1},\ldots,d_{n})=(c_{1}d_{1},\ldots,c_{n}d_{n}).$
In this section we start by enhancing such results to hold over any field $\mathbb{F}$}, i.e. we demonstrate that the proof of the Feng-Rao bounds hold regardless of underlying field. Our exposition closely follows that of [3, 25]. With such theory in hand and by applying Forney’s two duality lemmas [19] our results concerning multivariate polynomials follows as easy corollaries, but admittedly our exposition in the present section is more general than what is needed for that purpose as we want to support possible future work on generalizing our results in Section 2 to also hold over other algebraic structures than fields.
Consider a subspace $D\subseteq{\mathbb{F}}^{n}$ and define the support as
${\mbox{Supp}}(D)=\{i\in{\mathcal{I}}\mid c_{i}\neq 0{\mbox{ for some }}(c_{1},\ldots,c_{n})\in D\}.$
We are interested in establishing information on the size of ${\mbox{Supp}}(D)$ which we call $w_{H}(D)=\#{\mbox{Supp}}(D)$. Here, the subscript $H$ refers to the Hamming weight from coding theory.
Let $B=\{{\vec{b}}_{1},\ldots,{\vec{b}}_{n}\}$ be an ordered basis for ${\mathbb{F}}^{n}$. Define $L_{0}=\{\vec{0}\}$ and for $i=1,\ldots,n$ $L_{i}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{i}\}$. Further define $\bar{\rho}:{\mathbb{F}}^{n}\rightarrow\{0,1,\ldots,n\}$ by $\bar{\rho}(\vec{0})=0$ and for $\vec{c}\neq\vec{0}$ $\bar{\rho}(\vec{c})=i$ if $\vec{c}\in L_{i}\backslash L_{i-1}$ and $m(\vec{c})=\min\{m\mid\vec{c}\cdot\vec{b}_{m}\neq 0\}$. Here, $\cdot$ is the usual inner product.
Lemma 13.
Consider a subspace $D\subseteq{\mathbb{F}}^{n}$ of dimension $k\geq 1$. There exists a basis $\{\vec{c}_{1},\ldots,\vec{c}_{k}\}$ for $D$ with $\bar{\rho}(\vec{c}_{1}),\ldots,\bar{\rho}(\vec{c}_{k})$ being pairwise different. For such a basis it holds that $m(\vec{c}_{1}),\ldots,m(\vec{c}_{k})$ are also pairwise different. Both the $\bar{\rho}$-values and the $m$-values are invariants of $D$. If $\vec{c}\in D\backslash\{\vec{0}\}$ then $\bar{\rho}(\vec{c})$ and $m(\vec{c})$, respectively, belongs to the above mentioned sets, respectively.
Hence, the following definition is clear
$\bar{\rho}(D)=\{e_{1},\ldots,e_{k}\mid 1\leq e_{1}<\cdots<e_{k}\leq n,\forall e_{i}\exists\vec{c}\in D\backslash\{\vec{0}\}{\mbox{ with }}\vec{c}\in L_{e_{i}}\backslash L_{e_{i}-1}\}$
$m(D)=\{m_{1},\ldots,m_{k}\mid 1\leq m_{1}<\cdots<m_{k}\leq n,\forall m_{i}\exists\vec{c}\in D\backslash\{\vec{0}\}{\mbox{ with }}\vec{c}\in L^{\perp}_{m_{i}-1}\backslash L^{\perp}_{m_{i}}\}$
We next introduce the concept of one-way well-behaving pairs which is more general than the more well-known concepts of well-behaving pairs and weakly well-behaving pairs [32] allowing for a more powerful description [25]. Consider an additional ordered basis $B^{\prime}=\{\vec{b}^{\prime}_{1},\ldots\vec{b}^{\prime}_{n}\}$ (which may or may not be equal to $B$). Let ${\mathcal{I}}:=\{1,2,\ldots,n\}$. An ordered pair $(i,j)\in{\mathcal{I}}^{2}$ is said to be one-way well-behaving (OWB) if $\bar{\rho}(\vec{b}_{u}\ast\vec{b}^{\prime}_{j})<\bar{\rho}(\vec{b}_{i}\ast\vec{b}^{\prime}_{j})$ for $u<i$. Given bases $B,B^{\prime}$ as above consider for $l=1,2,\ldots,n$ and $i=1,2,\ldots,n$ the following sets
${\mbox{V}}_{l}$ $:=$ $\{i\in I\mid\bar{\rho}(\vec{b}_{i}\ast\vec{b}_{j}^{\prime})=l{\mbox{ for some }}\vec{b}_{j}^{\prime}\in B^{\prime}{\mbox{ with }}(i,j){\mbox{ OWB}}\}$ (6)
$\Lambda_{i}$ $:=$ $\{l\in I\mid\bar{\rho}(\vec{b}_{i}\ast\vec{b}_{j}^{\prime})=l{\mbox{ for some }}\vec{b}_{j}^{\prime}\in B^{\prime}{\mbox{ with }}(i,j){\mbox{ OWB}}\}$ (7)
Definition 14.
For $\{l_{1},\ldots,l_{k}\}\subseteq I$ and $\{i_{1},\ldots,i_{k}\}\subseteq I$ define
$\bar{\mu}(l_{1},\ldots l_{k})$ $:=$ $\#\left(\left(\cup_{s=1,\ldots,k}{\mbox{V}}_{l_{s}}\right)\cup\{l_{1},\ldots,l_{k}\}\right)$ (8)
$\bar{\sigma}(i_{1},\ldots i_{k})$ $:=$ $\#\left(\left(\cup_{s=1,\ldots,k}\Lambda_{i_{s}}\right)\cup\{i_{1},\ldots,i_{k}\}\right)$ (9)
Theorem 15.
Let $D\subseteq{\mathbb{F}}^{n}$. Then it holds that
$w_{H}(D)$ $\geq$ $\bar{\sigma}(\bar{\rho}(D))$ (10)
$w_{H}(D)$ $\geq$ $\bar{\mu}(m(D)).$ (11)
Proof.
By inspection the proofs in [3] and [25] do not use any assumption on the field under consideration.
To prove (10) we therefore repeat the proof of [3][Th. 10] which with our notation goes as follows. Write $e_{u}:=\bar{\rho}(\vec{c}_{u})$ for $u=1,\ldots k$ and observe that if $(e_{u},j)$ is OWB for some $j\in\{1,\ldots,n\}$ and $\bar{\rho}(\vec{b}_{e_{u}}\ast\vec{b}^{\prime}_{j})=\ell$ then by the very definition of OWB also $\bar{\rho}(\vec{c}_{u}\ast\vec{b}_{j}^{\prime})=\ell$. Hence,
$W:=\cup_{u=1}^{k}\{\vec{c}_{u}\ast\vec{b}^{\prime}_{j}\mid(a_{u},j){\mbox{ is OWB}}\}$
contains at least $\left(\cup_{u=1}^{k}\Lambda_{e_{u}}\right)$ linearly independent vectors. Consequently,
$W^{\prime}:=W\cup\{\vec{c}_{u}\ast(1,\ldots,1)\mid u=1,\ldots,k\}$
contains at least $\bar{\sigma}(e_{1},\ldots,e_{k})$ linearly independent vectors. From this we conclude
${\mbox{Supp}}D\geq\dim{\mbox{ Span}}_{\mathbb{F}}W^{\prime}=\bar{\sigma}(\bar{\rho}(D)).$
To prove (11) we repeat the proof of the last part of [25][Th. 1]. Write $\gamma:=\bar{\mu}(m_{1}:=m(\vec{c}_{1}),\ldots,m_{k}:=m(\vec{c}_{t}))$ and
$\begin{array}[]{l}\{i_{1},\ldots,i_{\gamma}\}:=\\ \cup_{s=1}^{k}\left(\{i\in{\mathcal{I}}\mid\exists\vec{b}^{\prime}_{j}\in{\mathcal{B}}^{\prime}{\mbox{ with }}\bar{\rho}(\vec{b}_{i}\ast\vec{b}^{\prime}_{j})=m_{s}{\mbox{ and }}(i,j){\mbox{ OWB}}\}\cup\{m_{s}\}\right)\end{array}$
where without of loss of generality we assume $i_{1}<\cdots<i_{\gamma}$. For $1\leq h\leq\gamma$ consider
$\vec{r}_{h}=\sum_{v=1}^{h}\alpha_{v}\vec{b}_{i_{v}},\,\,\alpha_{v}\in{\mathbb{F}},\,\,\alpha_{h}\neq 0$
and let $T$ be the vector-space consisting of all possible $\vec{r}_{h}$ as well as $\vec{0}$.
If $i_{h}\in\{m_{1},\ldots,m_{k}\}$ then by the very definition of the function $m$ we have $\vec{r}_{h}\ast\vec{c}_{h}\neq 0$ and consequently $\vec{r}_{h}\ast\vec{c}_{h}\neq\vec{0}$. If $i_{h}\notin\{m_{1},\ldots,m_{k}\}$ then there exists a $j$ and an $m_{u}$, $u\in\{1,\ldots,k\}$ such that $\bar{\rho}(\vec{b}_{i_{h}}\ast\vec{b}^{\prime}_{j})=m_{u}$ with $(i_{h},j)$ OWB. It follows that $(\vec{r}_{h}\ast\vec{b}^{\prime}_{j})\cdot\vec{c}_{u}\neq 0$ from which we deduce that $\vec{r}_{h}\ast\vec{c}_{u}\neq\vec{0}$. In conclusion for every non-zero element $\vec{r}_{h}$ of $T$ there exists a $\vec{c}\in D$ with $\vec{r}_{h}\ast\vec{c}\neq\vec{0}$. It follows that ${\mbox{Supp}}(D)\geq\dim T=\gamma$, and we are through. ∎
Remark 16.
For the purpose of the present paper the reader may think of $\mathcal{B}$} and $\mathcal{B}$$^{\prime}$ as the same basis. We note, that our results hold also in the general setting where $\mathcal{B}$$^{\prime}$ is any set, not necessarily a basis. In coding theory there are examples where the concept of one-way well-behaving gives better result than the symmetric concepts of well-behaving and weakly well-behaving pairs. We leave it for further research to possibly enhance the theory of the present paper to include structures over skew fields where possibly the one-way feature might be instrumental.
To address (II) at the level of linear algebra it is enough to combine (11) with Forney’s duality-lemmas from [19] which we now demonstrate. The concepts of puncturing and projection, respectively, are given as follows. Given a subspace $C\subseteq{\mathbb{F}}^{n}$ and $A\subseteq{\mathcal{I}}$ we write
$C_{A}:=\{\vec{c}=(c_{1},\ldots,c_{n})\in C\mid c_{i}=0{\mbox{ for all }}i\in{\mathcal{I}}\backslash A=:\bar{A}\}.$
For $\vec{c}=(c_{1},\ldots,c_{n})$ we define ${\mathcal{P}}_{A}(\vec{c})=(d_{1},\ldots,d_{n})$ where $d_{i}=c_{i}$ whenever $i\in A$ and $d_{i}=0$ otherwise from which we define
${\mathcal{P}}_{A}(C):=\{{\mathcal{P}}_{A}(\vec{c})\mid\vec{c}\in C\}.$
Forney’s duality lemmas [19][Lem. 1 and Lem. 2] read
$\dim C$ $=$ $\dim C_{\bar{A}}+\dim{\mathcal{P}}_{A}(C)$ (12)
$\#A$ $=$ $\dim{\mathcal{P}}_{A}(C)+\dim(C^{\perp})_{A}$ (13)
To ease the exposition we now recall the concept of relative generalized Hamming weights, but here described in the setting of arbitrary field (we already applied such notation in (4)). Given $C_{2}\subsetneq C_{1}\subseteq{\mathbb{F}}^{n}$ write $k_{2}=\dim C_{2}$, $k_{1}=\dim C_{1}$ and $t=k_{1}-k_{2}$. For $1\leq k\leq t$ we define
$M_{k}(C_{1},C_{2}):=\min\{\#{\mbox{Supp}}(D)\mid D\subseteq C_{1},D\cap C_{2}=\{\vec{0}\},\dim D=k\}$
and similarly for $C_{1}^{\perp}\subsetneq C_{2}^{\perp}$.
Observe, that in the particular case
$C_{2}$ $=$ ${\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{2}}\}$
$C_{1}$ $=$ ${\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{1}}\}$
by Theorem 15 we obtain
$M_{k}(C_{1},C_{2})$ $\geq$ $\min\{\bar{\sigma}(e_{1},\ldots,e_{k})\mid k_{2}<e_{1}<\cdots<e_{k}\leq k_{1}\}$ (14)
$M_{k}(C_{2}^{\perp},C_{1}^{\perp})$ $\geq$ $\min\{\bar{\mu}(e_{1},\ldots,e_{k})\mid k_{2}<e_{1}<\cdots<e_{k}\leq k_{1}\}.$ (15)
We are now ready to state our answer to (II) in the setting of linear algebra based on which in the next section we shall prove Theorem 5.
Theorem 17.
Let $1\leq k_{2}<k_{1}\leq n$ and consider $C_{2}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{2}}\}\subseteq C_{1}:={\mbox{Span}}_{\mathbb{F}}\{\vec{b}_{1},\ldots,\vec{b}_{k_{1}}\}$. Write $t=k_{1}-k_{2}$ and consider $1\leq k\leq t$. For all $A\subseteq{\mathcal{I}}$ with
$A<\min\{\bar{\mu}(i_{1},\ldots,i_{t-k+1})\mid k_{2}+1\leq i_{1}<\cdots<i_{t-k+1}\leq k_{1}\}$ (16)
there exist $k$ linearly independent vectors in $C_{1}\backslash C_{2}$ that are identically equal to 0 on $A$.
When $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ equals the right-hand side of (16) for some $A$ of size equal to this value there does not exist $k$ such linearly independent vectors.
Proof.
We prove that for all $A\subseteq{\mathcal{I}}$ with $A<M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ the result holds and that for some $A$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ it does not. Combining this with (15) (where we substitute $k$ with tk+1t-k+1) finalizes the proof. Given $A$ the number of linearly independent vectors being identically equal to 0 on $A$ is
$\dim(C_{1})_{\bar{A}}-\dim(C_{2})_{\bar{A}}$
$=$ $(k_{1}-\dim P_{A}(C_{1}))-(k_{2}-\dim P_{A}(C_{2}))$
$=$ $t-\left(\dim{{P}}_{A}(C_{1})-\dim P_{A}(C_{2})\right)$
$=$ $t-\left((\#A-\dim(C_{1}^{\perp})_{A})-(\#A-\dim(C_{2}^{\perp})_{A})\right)$
$=$ $t-\left(\dim(C_{2}^{\perp})_{A}-\dim(C_{1}^{\perp})_{A}\right)$
$=$ $t-\max\{\dim D\mid D\subseteq C_{2}^{\perp},D\cap C_{1}^{\perp}=\{\vec{0}\},{\mbox{Supp}}(D)\subseteq A\}$
where we used Forney’s two duality lemmas (12) and (13). By the very definition of relative generalized Hamming weights for any $A\subseteq{\mathcal{I}}$ of size strictly less than $M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ we have that
$\max\{\dim D\mid D\subseteq C_{2}^{\perp},D\cap C_{1}^{\perp}=\{\vec{0}\},{\mbox{Supp}}(D)\subseteq A\}$
is strictly smaller than tk+1t-k+1 and that for some $A\subseteq{\mathcal{I}}$ with $A=M_{t-k+1}(C_{2}^{\perp},C_{1}^{\perp})$ equality holds. This concludes the proof. ∎
4 Proof of Theorem 5
With the theory from Section 3 in place we are now ready to prove Theorem 5.
Proof of Theorem 5
If $\mathbb{F}$} is a finite field, say with $q$ elements, then consider $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$. Otherwise, let $q=\mu(M_{1},\ldots,M_{t})$ and choose finite sets $A_{1},\ldots,A_{m}\subseteq{\mathbb{F}}$ all of size at least $q$ in such a way that the $A$ in Theorem 5 satisfies $A\subseteq A_{1}\times\cdots\times A_{m}$. Write $a_{i}=\#A_{i}$, for $i=1,\ldots,m$, $n=a_{1}\cdots a_{m}$ and $\{P_{1},\ldots,P_{n}\}=A_{1}\times\cdots\times A_{m}$. Define $G_{i}:=\prod_{b\in A_{i}}(X_{i}-b)$, for $i=1,\ldots,m$ and consider the ideal $I=\langle G_{1},\ldots,G_{m}\rangle$. The footprint is
$\Delta_{\prec}(I)=\{X_{1}^{i_{1}}\cdots X_{m}^{i_{m}}\mid 0\leq i_{j}<a_{j},{\mbox{ for }}j=1,\ldots,m\}$
and by interpolation the map ${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n}$ given by ${\mbox{ev}}(F+I)=(F(P_{1}),\ldots,F(P_{n}))$ is surjective. Preimage and image are of the same size, and therefore $\{{\mbox{ev}}(N)\mid N\in\Delta_{\prec}(I)\}$ is a basis for ${\mathbb{F}}^{n}$ as a vector space over $\mathbb{F}$}. We enumerate the monomials of the footprint according to $\prec$ as $N_{1}\prec\cdots\prec N_{n}$ and write $\vec{b}_{i}={\mbox{ev}}(N_{i})$. From the identity ${\mbox{ev}}(F+I)\ast{\mbox{ev}}(G+I)={\mbox{ev}}(FG+I)$ we immediately see that $\mu(N_{i_{1}},\ldots,N_{i_{k}})\leq\bar{\mu}(i_{1},\ldots,i_{k})$ and therefore the first part of Theorem 5 follows from the first part of Theorem 17. In the remaining part of the proof we assume that $\mathbb{F}$} is the finite field ${\mathbb{F}}_{q}$. Using notation as in [9] consider $W\subseteq\Delta_{\prec}(I)$ and $E(W):={\mbox{Span}}_{\mathbb{F}_{q}}\{{\mbox{ev}}(M)\mid M\in W\}$ and $C(W):=E(W)^{\perp}$. By [9, Pro. 2.4] whenever $W$ is divisor closed it holds that $C(W)=E(W^{\perp})$ where
$W^{\perp}=\Delta_{\prec}(I)\backslash\{X_{1}^{q-1-e_{1}}\cdots X_{m}^{q-1-e_{m}}\mid X_{1}^{e_{1}}\cdots X_{m}^{e_{m}}\in W\}.$ (17)
Using the notation as in Theorem 5 in combination with that of Theorem 17 define
$C_{2}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{2}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec M_{1}\}\right)$
$C_{1}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{1}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq M_{t}\}\right)$
where of course ${\mbox{lm}}({\mathcal{L}}_{1}\backslash{\mathcal{L}}_{2})=\{M_{1},\ldots,M_{t}\}$. From (17) we obtain
$C_{2}^{\perp}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{2}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\preceq\dfrac{N}{M_{1}}\}\right)$
$C_{1}^{\perp}$ $=$ ${\mbox{ev}}\left({\mathcal{L}}_{1}^{\perp}:={\mbox{Span}}_{\mathbb{F}_{q}}\{M\mid M\in\Box,M\prec\dfrac{N}{M_{t}}\}\right)$
where $N=X_{1}^{q-1}\cdots X_{m}^{q-1}$. Clearly ${\mbox{lm}}({\mathcal{L}}^{\perp}_{2}\backslash{\mathcal{L}}_{1}^{\perp})=\left\{\dfrac{N}{M_{t}},\ldots,\dfrac{N}{M_{1}}\right\}$ and the last part of Theorem 5 follows from the last part of Theorem 17 in combination with the last part of Corollary 10.
$\Box$
Remark 18.
Inspecting the above proof of Theorem 5 one can fill in the details missing in Remark 11.
Alternative proof of Corollary 10
Using similar terminology as in the previous proof one can demonstrate that Corollary 10 is a consequence of (14). We leave the details for the reader.
$\Box$
5 Concluding remarks
In this paper we proposed a unified theory for common affine roots of general sets of multivariate polynomials over any field, except we did not treat multiplicity. For multivariate polynomials there are several concepts of multiplicity but in particular the one related to Hasse-derivatives seems to have gained interest in the literature [38, 34, 13, 39, 15, 26, 14]. We leave it for further research to investigate if the results in there can be further developed to possibly obtain sharper findings along the line of Theorem 5, Corollary 10 and Remark 11.
We further pose it as a research problem to enhance our findings to hold for integral domains, skew fields and possibly non-commutative integer rings generalizing results in [7, 35]. Also we believe that the material of Section 3 should have implications for the study of particular types of point-sets such as roots of absolutely irreducible polynomials etc. Similarly it would be interesting to see if some of the methods in the present paper can be applied to treat projective points beyond what is already in the litterature e.g. in [11, 5, 33]
We note that like many other recent results on roots of sets of multivariate polynomials also the present paper is inspired by work in coding theory and cryptography. For this particular paper the inspiration comes from work on ramp secret sharing schemes [22].
The author sincerely thanks Anurag Bishnoi for a fruitful email correspondence during the making of this paper.